# ImpMIA: Leveraging Implicit Bias for Membership Inference Attack under Realistic Scenarios

### v1

@@ B.5 Weight Decay Influence
Without weight decay 0.70 1.69 0.99
$10^{-6}$ 0.71 1.73 0.93
$10^{-5}$ 0.70 2.04 1.09
$\bm{10^{-4}}$ 0.71 1.48 0.90
B.6 Influence of Non-Member Ratio on Evaluation
Our assumption-removal analysis also showed the influence of candidate pool size on model performance. In this setting, we lowered the member ratio by adding more non-members to the superset (increasing from 25K to 55K non-members). While our method does not rely on any specific member-to-non-member ratio, adding more non-members can affect results in two ways: (i) it increases the number of samples, which may make it harder to classify members, and (ii) it directly affects the evaluation metric, since FPR is defined relative to the total number of non-member samples. The overall performance reflects the interaction of these two factors. On the one hand, a larger pool introduces more distractors, potentially reducing accuracy. On the other hand, at fixed FPR thresholds (e.g., 0.01%), a larger pool allows more absolute mistakes. For instance, with 25K non-members, 0.01% FPR corresponds to only two false positives, whereas with 55K non-members, it allows up to five. Since coefficients are not uniformly distributed, permitting more mistakes can yield a nonlinear gain in TPR. In practice, we observed that at 0.01% FPR, our method performed better with 55K non-members than with 25K. This suggests that the positive effect of the evaluation metric outweighs the negative impact of additional distractors. Importantly, at 0% FPR—where the metric is unaffected by pool size—the results remained stable, as expected.
@@ C.1 Black-Box Baselines
For black-box comparisons, we used the official RMIA implementation, which also includes code for LiRA, Attack-P, and Attack-R. We ran the full-power RMIA variant described by Zarifzadeh et al. (2023), which trains 256 reference models per dataset, and used the same framework to evaluate the other black-box baselines. To ensure consistency, we adapted the code to match the training configuration of Cohen & Giryes (2024) (SIF): ResNet-18 backbone, inputs normalized to $[0,1]$, and standard augmentations (random crop and horizontal flip).
RMIA requires dataset-specific configurations, which we followed exactly as provided in their code and paper for CIFAR-10, CIFAR-100, and CINIC-10. All experiments were therefore run with the recommended hyperparameters for each dataset. We emphasize that RMIA, despite reporting state-of-the-art results in its original paper, is highly sensitive to training configurations—especially at very low false-positive rates. Even minor mismatches in normalization, architecture, optimizer, or learning-rate schedule caused severe degradation in our experiments.
@@ C.2 White-Box Baselines
Nasr et al. (Nasr et al., 2019) showed that the most informative white-box membership signal is the magnitude of per-sample gradients, since stochastic gradient descent drives member gradients toward zero. Building on this, we evaluate two baselines: (i) a loss-gradient score based on the norm $\|\nabla_{\theta}L(f(x),y)\|$, and (ii) a margin-gradient score based on the gradient $\|\nabla_{\theta}[f_{y}(x)-\max_{j\neq y}f_{j}(x)]\|$. For each sample (and its horizontal flip), we compute per-layer gradient norms, convert them into rank-based scores (higher rank = smaller norm), and average across layers and augmentations. The loss-gradient baseline directly instantiates Nasr’s observation, while the margin-gradient baseline adapts it to sensitivity around the decision boundary, making it conceptually closer to ImpMIA.
@@
The self-influence function (SIF) attack of Cohen & Giryes (2024) achieved state-of-the-art white-box membership inference in their paper, particularly under strong augmentations where gradient-norm methods fail. SIF measures a sample’s effect on its own loss by approximating the influence function via recursive Hessian–vector products. Membership is inferred by whether a correctly classified sample’s score lies close to zero, since training points tend to cluster tightly around this value while non-members yield more extreme values.
@@ C.3 Additional Results
(c) CINIC-10
Figure S2: TPR–FPR plots for the no-assumption combined setting. These curves illustrate attack performance when the attacker faces realistic uncertainty: (i) training hyperparameters are unknown, (ii) the candidate pool mixes in- and out-of-distribution samples (distribution shift), and (iii) the fraction of members is unknown. The plots complement the main text by showing the full ROC behavior, especially in the low-FPR regime.
Table T4: Membership Inference Results (All Assumptions). Performance across all datasets under the standard assumption-rich setting where training hyperparameters, data distribution, and member ratio are known. Metrics reported are TPR (%) at 0.00% and 0.01% FPR, with AUC included for completeness.
@@
GradNorm – loss 0.61 0.01 0.00 0.83 0.01 0.00 0.74 0.00 0.00
GradNorm – margin 0.59 0.00 0.00 0.72 0.04 0.01 0.69 0.01 0.00
AdaSIF 0.62 0.01 0.00 0.87 0.03 0.01 0.75 0.00 0.00
@@
Table T5: Membership Inference Results (Unknown Member Fraction). CINIC-10 results when the attacker does not know the proportion of training members in the candidate pool. Metrics reported are TPR (%) at 0.00% and 0.01% FPR, plus AUC.
@@
Table T6: Membership Inference Results (Different Distribution). Performance across all datasets when the candidate pool mixes in-distribution and out-of-distribution data. Metrics reported are TPR (%) at 0.00% and 0.01% FPR, with AUC included for completeness.
@@
Table T7: Membership Inference Results (Unknown Training Configuration). Performance across all datasets when the attacker does not know the target model’s training hyperparameters. Metrics reported are TPR (%) at 0.00% and 0.01% FPR, with AUC included for completeness.
@@
Table T8: LiRA and RMIA: Offline vs. Online under No-Assumptions. Results for LiRA and RMIA across all datasets in the realistic no-assumptions setting. The offline variant trains reference models independently of the candidate superset, while the online variant trains reference models directly on subsets of the superset, closer to the evaluation setup. Metrics reported are TPR (%) at 0.00% and 0.01% FPR, with AUC included for completeness.
@@
RMIA (offline) 0.79 0.26 0.00 0.97 7.35 1.67 0.86 0.11 0.02
Table T9: Membership Inference Results (No Assumptions). Performance across all datasets under the realistic no-assumptions setting. Metrics reported are TPR (%) at 0.00% and 0.01% FPR, with standard error included. This table shows the same results as in Table 1, but with standard error values reported.
@@
RMIA 0.19 $\pm$ 0.04 0.01 $\pm$ 0.00 6.73 $\pm$ 0.84 1.22 $\pm$ 0.45 0.15 $\pm$ 0.02 0.03 $\pm$ 0.00
GradNorm–loss 0.11 $\pm$ 0.01 0.01 $\pm$ 0.00 0.10 $\pm$ 0.02 0.04 $\pm$ 0.01 0.09 $\pm$ 0.02 0.01 $\pm$ 0.00
GradNorm–margin 0.02 $\pm$ 0.01 0.00 $\pm$ 0.00 0.02 $\pm$ 0.01 0.01 $\pm$ 0.00 0.03 $\pm$ 0.00 0.01 $\pm$ 0.01
AdaSIF 0.05 $\pm$ 0.01 0.00 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00
@@
Table T10: Assumptions Influence (CINIC-10, @0.01% FPR). Each entry is TPR (%) $\pm$ se. This table shows the same results as in Table 2, but with standard error values reported.
Method
Standard
Setting
Unknown
Config
Different
Distribution
Unknown
Member Ratio
No
Assumptions
@@
Table T11: Assumptions Influence (CINIC-10, @0.00% FPR). Each entry is TPR (%) $\pm$ se. This table shows the same results as in Table 2, but with standard error values reported.
Method
Standard
Setting
Unknown
Config
Different
Distribution
Unknown
Member Ratio
No
Assumptions
Attack-R 1.81 $\pm$ 0.55 0.37 $\pm$ 0.26 1.42 $\pm$ 0.80 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
@@ Appendix D Binary case of the KKT equations and Implicit Bias with Weight Decay
For completeness, we also present the implicit bias of neural networks for the binary classification case, as appeared in Lyu & Li (2019); Ji & Telgarsky (2020); Haim et al. (2022), and its extension to training with weight decay that was previously considered in (Buzaglo et al., 2023).
@@
This relation shows that the trained weights again lie in the span of per-sample gradients, with coefficients $\{\ell_{i}^{\prime}\}$ determined by the derivative of the loss. Importantly, equation 8 is structurally equivalent to the stationarity condition in the KKT system of the max-margin problem: in both formulations, the parameters are expressed as a linear combination of margin-gradient directions. Furthermore, if $\ell$ is the logistic loss function, then the coefficients $\{\ell_{i}^{\prime}\}$ of this linear combination are non-negative.
@@ Appendix E Limitations of Average-Case Metrics
This mismatch is evident in our results. Table 1 reports performance of several attacks in the ”No Assumptions” setting. On CIFAR-10, both our attack and the GradNorm baseline achieve AUC $\approx 0.81$. However, this similarity is misleading: GradNorm leverages model confidence, and its apparent effectiveness comes from the fact that non-members in this scenario has low confidence. By contrast, as shown by Haim et al. (2022), the truly memorized samples are those near the decision boundary. Our attack explicitly targets such near-margin points, yielding much higher TPR at very low FPR (e.g., $2.76\%$ vs. $0.11\%$ at $0.01\%$ FPR). This gap illustrates why average-case metrics like AUC are problematic: they make GradNorm appear competitive, while in reality it fails where privacy evaluation matters most—the ultra–low-FPR regime

### v3

@@ B.6 Weight Decay Influence
Without weight decay 0.81 2.93 1.62
$\bm{10^{-4}}$ 0.81 2.76 1.41
B.7 Influence of Pre-Filtering
The pre-filtering step aims to both stabilize the optimization by eliminating samples that are unlikely to be members (and therefore may add noise) and speed up optimization by having fewer samples. On the other hand, it may discard data points that are hard to learn even though they were in the training data, and therefore miss member samples. Since any attack is likely to miss some members, this is a trade-off that can be chosen in practice.
We analyze the impact of the pre-filtering step on our method’s performance in the CIFAR-10 No-Auxiliary-Knowledge setting. On CIFAR-10, pre-filtering removes between 4–25% of the examples, depending on the seed and scenario. As shown in Tab. T6, using pre-filtering leads to a slight increase in AUC (from 0.78 to 0.81), as well as improvements in TPR at 0.01% FPR (from 2.28 to 2.76) and at 0.0% FPR (from 1.15 to 1.41). Overall, the pre-filtering step does not appear to have a significant influence on performance, but it does speed up optimization by reducing the number of points being optimized
Variant AUC @ 0.01% FPR @ 0.0% FPR
No pre-filtering 0.78 2.28 1.15
With pre-filtering 0.81 2.76 1.41
Table T6: Effect of pre-filtering on ImpMIA (CIFAR-10, no auxiliary knowledge). Comparison of AUC and TPR at very low FPR with and without the pre-filtering step. We report AUC and TPR at fixed false-positive rates (FPR = 0.01% and 0.0%).
B.8 Running Time
The runtime of ImpMIA ranges between 12–16 hours on a single H200 GPU, depending on the size of the candidate pool (50K–80K samples), and can be reduced to approximately 2 hours using 8 GPUs. The black-box reference-model attacks in their strongest setting, used in this paper as well, require training 256 reference models. In the Full Auxiliary-Knowledge setting (400 epochs with a candidate pool of size 50K), this results in a total runtime of approximately 48 hours on a single GPU, which is roughly $4\times$ slower than ImpMIA. Notably, the runtime of black-box reference-model attacks is highly influenced by the number of epochs assumed by the attacker to mimic the target model’s training (which can be both lower or much higher than 400). In contrast, the runtime of ImpMIA remains largely unchanged, as it requires no reference model training.
@@ C.1 Black-Box Baselines
For black-box comparisons, we used the official RMIA implementation, which also includes code for LiRA, Attack-P, and Attack-R. We ran the full-power RMIA variant described by Zarifzadeh et al. (2023), which trains 256 reference models per dataset, and used the same framework to evaluate the other black-box baselines. To ensure consistency, we adapted the code to match the training configuration of Cohen and Giryes (2024) (SIF): ResNet-18 backbone, inputs normalized to $[0,1]$, and standard augmentations (random crop and horizontal flip).
RMIA requires dataset-specific configurations, which we followed exactly as provided in their code and paper for CIFAR-10, CIFAR-100, and CINIC-10. All experiments were therefore run with the recommended hyperparameters for each dataset. We emphasize that RMIA, despite reporting state-of-the-art results in its original paper, is highly sensitive to training configurations, especially at very low false-positive rates. Even minor mismatches in normalization, architecture, optimizer, or learning-rate schedule caused severe degradation in our experiments.
@@ C.2 White-Box Baselines
(Nasr et al., 2019) showed that the most informative white-box membership signal is the magnitude of per-sample gradients, since stochastic gradient descent drives member gradients toward zero. Building on this, we evaluate two baselines: (i) a loss-gradient score based on the gradients of the loss $\|\nabla_{\theta}L(f(x),y)\|$, and (ii) a margin-gradient score based on the gradients of the margin $\|\nabla_{\theta}[f_{y}(x)-\max_{j\neq y}f_{j}(x)]\|$. For each sample (and its horizontal flip), we compute per-layer gradient norms, convert them into rank-based scores (higher rank = smaller norm), and average across layers and augmentations. The loss-gradient baseline directly instantiates (Nasr et al., 2019) observation, while the margin-gradient baseline adapts it to sensitivity around the margin, making it conceptually closer to ImpMIA.
@@
The self-influence function (SIF) attack of Cohen and Giryes (2024) achieved state-of-the-art white-box membership inference in their paper, particularly under strong augmentations where gradient-norm methods fail. SIF measures a sample’s effect on its own loss by approximating the influence function via recursive Hessian–vector products. Membership is inferred by whether a correctly classified sample’s score lies close to zero, since training points tend to cluster tightly around this value while non-members yield more extreme values.
@@ C.3 Additional Results
Figure S2: TPR–FPR plots for the No-Auxiliary-Knowledge setting. These curves illustrate attack performance when the attacker faces realistic uncertainty: (i) training hyperparameters are unknown, (ii) the candidate pool mixes in- and out-of-distribution samples (distribution shift), and (iii) the fraction of members is unknown. The plots complement the main text by showing the full ROC behavior, especially in the low-FPR regime.
@@
GradNorm – loss 0.61 0.01 0.0 0.83 0.01 0.0 0.74 0.0 0.0
GradNorm – margin 0.59 0.0 0.0 0.72 0.04 0.01 0.69 0.01 0.0
AdaSIF 0.62 0.01 0.0 0.87 0.03 0.01 0.75 0.0 0.0
@@
Table T7: Membership Inference results (full auxiliary knowledge). Performance across all datasets under the assumption-rich setting where training hyperparameters, data distribution, and member ratio are known. Metrics reported are TPR (%) at 0.0% and 0.01% FPR, with AUC included for completeness.
@@
Table T8: Membership Inference results (unknown fraction of members). CINIC-10 results when the attacker does not know the proportion of training members in the candidate pool. Metrics reported are TPR (%) at 0.0% and 0.01% FPR, plus AUC.
@@
Table T9: Membership Inference results (different data distribution). Performance across all datasets when the candidate pool mixes in-distribution and out-of-distribution data. Metrics reported are TPR (%) at 0.0% and 0.01% FPR, with AUC included for completeness.
@@
Table T10: Membership Inference results (unknown training configuration). Performance across all datasets when the attacker does not know the target model’s training hyperparameters. Metrics reported are TPR (%) at 0.0% and 0.01% FPR, with AUC included for completeness.
@@
RMIA (offline) 0.79 0.26 0.0 0.97 7.35 1.67 0.86 0.11 0.02
Table T11: LiRA and RMIA: offline vs. online (no auxiliary knowledge). Results for LiRA and RMIA across all datasets in the realistic no-auxiliary-knowledge setting. The offline variant trains reference models independently of the candidate superset, while the online variant trains reference models directly on subsets of the superset, closer to the evaluation setup. Metrics reported are TPR (%) at 0.0% and 0.01% FPR, with AUC included for completeness.
@@
RMIA 0.19 $\pm$ 0.04 0.01 $\pm$ 0.0 6.73 $\pm$ 0.84 1.22 $\pm$ 0.45 0.15 $\pm$ 0.02 0.03 $\pm$ 0.0
GradNorm–loss 0.11 $\pm$ 0.01 0.01 $\pm$ 0.0 0.10 $\pm$ 0.02 0.04 $\pm$ 0.01 0.09 $\pm$ 0.02 0.01 $\pm$ 0.0
GradNorm–margin 0.02 $\pm$ 0.01 0.0 $\pm$ 0.0 0.02 $\pm$ 0.01 0.01 $\pm$ 0.0 0.03 $\pm$ 0.0 0.01 $\pm$ 0.01
AdaSIF 0.05 $\pm$ 0.01 0.0 $\pm$ 0.0 0.01 $\pm$ 0.0 0.0 $\pm$ 0.0 0.01 $\pm$ 0.0 0.0 $\pm$ 0.0
@@
Table T12: Membership Inference Results (no auxiliary knowledge). Performance across all datasets under the realistic no-assumptions setting. Metrics reported are TPR (%) at 0.0% and 0.01% FPR, with standard error included. This table shows the same results as in Tab. 1, but with standard error values reported.
Method Full Auxiliary-Knowledge Unknown Training Config. Different Distribution Unknown Fraction of Members NO Auxiliary-Knowledge
@@
Table T13: Assumptions influence (CINIC-10, @0.01% FPR). Each entry is TPR (%) $\pm$ se. This table shows the same results as in Tab. 3, but with standard error values reported.
Method Full Auxiliary-Knowledge Unknown Training Config. Different Distribution Unknown Fraction of Members NO Auxiliary-Knowledge
Attack-R 1.81 $\pm$ 0.55 0.37 $\pm$ 0.26 1.42 $\pm$ 0.80 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0
@@
Table T14: Assumptions influence (CINIC-10, @0.0% FPR). Each entry is TPR (%) $\pm$ se. This table shows the same results as in Tab. 3, but with standard error values reported.
@@ Appendix D Binary case of the KKT equations and Implicit Bias with Weight Decay
For completeness, we also present the implicit bias of neural networks for the binary classification case, as appeared in Lyu and Li (2019); Ji and Telgarsky (2020); Haim et al. (2022); Smorodinsky et al. (2024), and its extension to training with weight decay that was previously considered in (Buzaglo et al., 2023).
@@
This relation shows that the trained weights again lie in the span of per-sample gradients, with coefficients $\{\ell_{i}^{\prime}\}$ determined by the derivative of the loss. Importantly, (10) is structurally equivalent to the stationarity condition in the KKT system of the max-margin problem: in both formulations, the parameters are expressed as a linear combination of margin-gradient directions. Furthermore, if $\ell$ is the logistic loss function, then the coefficients $\{\ell_{i}^{\prime}\}$ of this linear combination are non-negative.
@@ Appendix E Limitations of Average-Case Metrics
This mismatch is evident in our results. Table 1 reports performance of several attacks in the No-Auxiliary-Knowledge setting. On CIFAR-10, both our attack and the GradNorm baseline achieve AUC $\approx 0.81$. However, this similarity is misleading: GradNorm leverages model confidence, and its apparent effectiveness comes from the fact that non-members in this scenario have low confidence. By contrast, as shown by Haim et al. (2022), the truly memorized samples are those near the decision boundary. Our attack explicitly targets such near-margin points, yielding much higher TPR at very low FPR (e.g., $2.76\%$ vs. $0.11\%$ at $0.01\%$ FPR). This gap illustrates why average-case metrics like AUC are problematic: they make GradNorm appear competitive, while in reality it fails where privacy evaluation matters most, the low-FPR regime